# Synthetic Data Generation in Cybersecurity: A Comparative Analysis

### v1

@@ VIII Future Work
    Fidelity vs Privacy: Enhancing the fidelity and utility of synthetic data often poses privacy risks. Thus, it is imperative to identify an optimized balance between fidelity and privacy, particularly concerning cybersecurity datasets. This endeavor requires careful consideration to safeguard individual privacy while maintaining the data’s utility for cybersecurity purposes.
@@
IX Conclusion
Our comparative study systematically explores the effectiveness of synthetic tabular data generation techniques within cybersecurity network traffic, covering both daily and malicious traffic scenarios. Through a detailed evaluation of several synthetic data generation methods across two significant datasets, we have revealed critical insights into their performances and the inherent trade-offs they entail. SMOTE and Cluster Centroids demonstrated substantial data similarity and class balance but suffer from the drawback of producing redundant, non-novel values. Conversely, CTGAN and TVAE outperformed the other methods, although they do not inherently ensure balanced class distribution. CastGAN emerges as a strong contender by maintaining internal feature dependencies, offering a more balanced approach. The insights from this study provide essential guidance for selecting synthetic data generation techniques based on specific cybersecurity requirements. This research significantly impacts the field by highlighting the strengths and limitations of each method, offering a clear framework for optimizing synthetic data generation in cybersecurity applications. The source code of this experiment is available at our https://github.com/AdenRajput/Comparative_Analysis.git.
X Acknowledgements

### v2

@@ VII Future Work
    Fidelity vs Privacy: Enhancing the fidelity and utility of synthetic data often poses privacy risks. Thus, it is imperative to identify an optimized balance between fidelity and privacy, particularly concerning cybersecurity network traffic datasets. This endeavor requires careful consideration to safeguard individual privacy while maintaining the data’s utility.
@@
VIII Conclusion
This study systematically evaluates synthetic network traffic data generation methods, comparing statistical and AI-based approaches. Our findings reveal that while statistical methods effectively handle class imbalance, they lack the ability to preserve complex feature dependencies. GAN-based models, particularly CTGAN and CopulaGAN, offer the best trade-off between fidelity and utility, making them preferable for generating realistic synthetic network traffic data. Diffusion models, despite their recent advances, exhibit significant computational challenges when applied to large-scale tabular datasets. The study highlights the need for hybrid approaches integrating statistical and AI techniques to improve class balance while maintaining data realism. Future research should explore optimizing generative AI loss functions, benchmarking real versus AI-generated class balance, and addressing privacy concerns while enhancing synthetic data fidelity.
IX Acknowledgements
@@
Appendix A Complete Graph Results for NSL-KDD Dataset
Figure 5: Probability Distribution Comparison for NSL KDD (Part 1/18)
Figure 6: Probability Distribution Comparison for NSL KDD (Part 2/18)
Figure 7: Probability Distribution Comparison for NSL KDD (Part 3/18)
Figure 8: Probability Distribution Comparison for NSL KDD (Part 4/18)
Figure 9: Probability Distribution Comparison for NSL KDD (Part 5/18)
Figure 10: Probability Distribution Comparison for NSL KDD (Part 6/18)
Figure 11: Probability Distribution Comparison for NSL KDD (Part 7/18)
Figure 12: Probability Distribution Comparison for NSL KDD (Part 8/18)
Figure 13: Probability Distribution Comparison for NSL KDD (Part 9/18)
Figure 14: Probability Distribution Comparison for NSL KDD (Part 10/18)
Figure 15: Probability Distribution Comparison for NSL KDD (Part 11/18)
Figure 16: Probability Distribution Comparison for NSL KDD (Part 12/18)
Figure 17: Probability Distribution Comparison for NSL KDD (Part 13/18)
Figure 18: Probability Distribution Comparison for NSL KDD (Part 14/18)
Figure 19: Probability Distribution Comparison for NSL KDD (Part 15/18)
Figure 20: Probability Distribution Comparison for NSL KDD (Part 16/18)
Figure 21: Probability Distribution Comparison for NSL KDD (Part 17/18)
Figure 22: Probability Distribution Comparison for NSL KDD (Part 18/18)
Appendix B Complete Graph Results for CIC-IDS2017 Dataset
Figure 23: Probability Distribution Comparison for CIC-IDS2017 (Part 1/7)
Figure 24: Probability Distribution Comparison for CIC-IDS2017 (Part 2/7)
Figure 25: Probability Distribution Comparison for CIC-IDS2017 (Part 3/7)
Figure 26: Probability Distribution Comparison for CIC-IDS2017 (Part 4/7)
Figure 27: Probability Distribution Comparison for CIC-IDS2017 (Part 5/7)
Figure 28: Probability Distribution Comparison for CIC-IDS2017 (Part 6/7)
Figure 29: Probability Distribution Comparison for CIC-IDS2017 (Part 7/7)
Figure 30: Correlation Heatmap for CIC-IDS2017 (Part 1/2)
Figure 31: Correlation Heatmap for CIC-IDS2017 (Part 2/2)